# LLaVA Needs More Knowledge: Retrieval Augmented Natural Language Generation with Knowledge Graph for Explaining Thoracic Pathologies

### v1

@@ 5 Results and Discussions
5.1 Comparison with Other Methods
In this study, we evaluated the performance of our proposed KG-LLaVA framework against several well-established methods, including RATCHET [5], TieNet [28], and DPT [14], using the MIMIC-NLE [14] dataset. The results, as summarized in Table 1, clearly demonstrate that KG-LLaVA outperforms the previous methods across a range of evaluation metrics.
@@ 5.1 Comparison with Other Methods
Notably, while KG-LLaVA slightly outperforms RATCHET [5] in METEOR (15.1 vs. 14.1), the overall superior performance of KG-LLaVA across the other metrics underscores the strength of incorporating KG-RAG module into a vision-language framework. These results suggest that KG-LLaVA has the potential to set a new benchmark for generating NLEs in medical imaging tasks, particularly for diagnosing thoracic pathologies.
While [25] methods focused on optimizing model parameters, our approach leveraged the KG-RAG module and effective use of LoRA for fine-tuning, achieving superior performance without compromising the model’s complexity or parameter efficiency. This makes KG-LLaVA not only the best-performing model but also a robust and scalable solution for medical NLE generation.
@@
Table 3: Comparative analysis of the performance of Med-XPT, and KG-LLaVA across different RAG methods and without any RAG. The table includes results for NLG metrics such as BLEU-4 (B4), METEOR, ROUGE-L (R-L), and CIDEr. The "-" row shows results without any RAG integration, the "NLE" row represents results with natural language explanation-based RAG, and the "KG" row reflects the performance when the knowledge graph retrieval module is used.
5.2 Comparison of different LLMs
@@
5.3 Impact of Different RAG Methods
@@
Finally, we conducted a detailed comparison of the two frameworks—Med-XPT, and KG-LLaVA—across various configurations: without any Retrieval Augmented Generation (RAG), with standard NLE, and with our proposed KG Retrieval module. The results, as shown in Table 3, illustrate the impact of different RAG methods on the performance of these frameworks in generating accurate and contextually rich NLEs for thoracic pathologies. In the NLE configuration, where standard NLEs are generated without KG enhancement, both Med-XPT and KG-LLaVA exhibit strong performance, with KG-LLaVA slightly leading in most metrics. This indicates that while both frameworks leverage their respective architectures effectively, the pre-training knowledge embedded in KG-LLaVA likely contributes to its superior performance.
@@
Despite the significantly smaller size and less extensive training of the GPT-2 language model compared to LLaVA, Med-XPT—which utilizes GPT-2—achieves competitive performance, particularly when augmented with the KG and NLE-based RAG methods. This suggests that the incorporation of additional domain-specific knowledge can effectively compensate for the limitations of smaller models, enabling them to generate high-quality explanations.
@@
Notably, when comparing Med-XPT to the DPT [14] framework, which also uses GPT-2 but without the benefit of these advanced RAG methods, we observe a substantial improvement in performance. DPT’s lower scores across the board highlight the critical role that the KG and NLE-based RAG play in enhancing the model’s explanatory capabilities. This further emphasizes that even smaller models, when equipped with the right augmentation techniques, can perform on par with larger, more complex models like LLaVA.
@@
5.4 Effect of the Number of Retrieved Knowledge Graph Triplets (K) on NLE
We investigate the effect of the number of retrieved knowledge graph triplets (K) on NLE. As shown in Table 4, as K increases from 1 to 7, the model demonstrates consistent performance across BLEU-4, METEOR, and ROUGE-L, with the highest CIDEr score observed at K=7. This suggests that retrieving more triplets enhances the richness and relevance of the generated explanations, particularly when the complexity of the input increases. However, the slight fluctuations in other metrics indicate that there is an optimal balance to be struck between the amount of retrieved knowledge and its utility in generating precise explanations.
5.5 Comparison of Uni-modal and Cross-modal Retrieval
Table 5 shows the comparison of uni-modal retrieval and cross-modal retrieval (ours). The Cross-modal retrieval method demonstrates a clear advantage over the Uni-modal approach across all evaluation metrics, with a particularly notable improvement in CIDEr, where the score increases from 49.9 to 62.2. This substantial gain highlights the effectiveness of integrating both visual and textual modalities in the retrieval process, allowing the model to generate more contextually relevant and accurate Natural Language Explanations (NLEs). The Cross-modal approach enhances the model’s ability to interpret complex medical images, leading to higher-quality explanations that are more aligned with clinical expectations.
@@ 5.5 Comparison of Uni-modal and Cross-modal Retrieval
Table 4: Effect of the Number of Retrieved Knowledge Graph Triplets (K) on NLE. We present the impact of varying the number of retrieved knowledge graph triplets (K) on the performance of the KG-LLaVA model in generating Natural Language Explanations (NLEs). The evaluation metrics include BLEU-4 (B4), METEOR, ROUGE-L (R-L), and CIDEr.
@@
Table 5: Comparison of Uni-modal and Cross-modal Retrieval on NLE Performance. We compare the performance of the KG-LLaVA model when using Uni-modal versus Cross-modal retrieval methods for generating Natural Language Explanations (NLEs).
@@
5.6 Qualitative Results
6 Conclusion
In this paper, we introduced a novel approach for generating NLEs for thoracic pathologies by integrating the KG-RAG module into vision-language models. Our KG-RAG framework effectively enhances the accuracy and contextual relevance of NLEs by incorporating domain-specific knowledge. Evaluated across three distinct frameworks—KG-LLaVA, Med-XPT, and Bio-LLaVA—our method consistently outperformed established models like RATCHET, TieNet, and DPT on the MIMIC-NLE dataset, highlighting the robustness and versatility of the KG-RAG approach.

### v2

@@ 5 Results and Discussions
Comparison with Other Methods
In this study, we evaluated the performance of our proposed KG-LLaVA framework against several well-established methods, including RATCHET (Hou et al. 2021), TieNet (Wang et al. 2018), and DPT (Kayser et al. 2022), using the MIMIC-NLE (Kayser et al. 2022) dataset. The results, as summarized in Table 1, clearly demonstrate that KG-LLaVA outperforms the previous methods across a range of evaluation metrics.
@@
Notably, while KG-LLaVA slightly outperforms RATCHET (Hou et al. 2021) in METEOR (15.1 vs. 14.1), the overall superior performance of KG-LLaVA across the other metrics underscores the strength of incorporating KG-RAG module into a vision-language framework. These results suggest that KG-LLaVA has the potential to set a new benchmark for generating NLEs in medical imaging tasks, particularly for diagnosing thoracic pathologies.
While (Rio-Torto, Cardoso, and Teixeira 2024) methods focused on optimizing model parameters, our approach leveraged the KG-RAG module and effective use of LoRA for fine-tuning, achieving superior performance without compromising the model’s complexity or parameter efficiency. This makes KG-LLaVA not only the best-performing model but also a robust and scalable solution for medical NLE generation.
@@
Table 3: Comparative analysis of the performance of Med-XPT, and KG-LLaVA across different RAG methods and without any RAG. The table includes results for NLG metrics such as BLEU-4 (B4), METEOR, ROUGE-L (R-L), and CIDEr. The ”-” row shows results without any RAG integration, the ”NLE” row represents results with natural language explanation-based RAG, and the ”KG” row reflects the performance when the knowledge graph retrieval module is used.
Comparison of different LLMs
@@
Impact of Different RAG Methods
@@
We conducted a detailed comparison of the two frameworks—Med-XPT and KG-LLaVA—across various configurations: without any Retrieval Augmented Generation (RAG), with standard NLE, and with our proposed KG Retrieval module. The results, as shown in Table 3, illustrate the impact of different RAG methods on the performance of these frameworks in generating accurate and contextually rich NLEs for thoracic pathologies.
In the NLE configuration, where standard NLEs are generated without KG enhancement, both Med-XPT and KG-LLaVA exhibit strong performance, with KG-LLaVA slightly leading in most metrics. This indicates that while both frameworks leverage their respective architectures effectively, the pre-training knowledge embedded in KG-LLaVA likely contributes to its superior performance.
Additional discussion is provided in Appendix C of the supplementary material.
Role of Instructional Prompts
An ablation study on the role of instructional prompts is available in Appendix D.
Qualitative Results
Analysis of Failure Case
A detailed discussion of failure cases is provided in Appendix E.
Limitation
Our framework requires a KG model like RadGraph for constructing triplets. For new modalities, practitioners may need to adapt or train alternative graph-processing models. While our design is intended for research purposes, further clinical validation is necessary before real-world deployment.
6 Conclusion
In this paper, we introduced a novel approach for generating NLEs for thoracic pathologies by integrating the KG-RAG module into vision-language models. Our KG-RAG effectively enhances the accuracy and contextual relevance of NLEs by incorporating domain-specific knowledge. Evaluated across three distinct frameworks—KG-LLaVA, Med-XPT, and Bio-LLaVA—our method consistently outperformed established models like RATCHET, TieNet, and DPT on the MIMIC-NLE dataset, highlighting the robustness and versatility of the KG-RAG approach.
Acknowledgments
This work was supported in part by the Institute of Information and Communications Technology Planning and Evaluation (IITP) Grant funded by the Korea Government (MSIT) under Grant 2022-0-00078 (Explainable Logical Reasoning for Medical Knowledge Generation), Grant IITP-2024-RS-2023-00258649, Grant RS-2022-00155911, Grant 2021-0-02068 and by the National Research Foundation of Korea (NRF) Grant funded by the Korea Government (MSIT) under Grant RS-2024-00334321.
Supplementary Material
LLaVA Needs More Knowledge: Retrieval Augmented Natural Language Generation with Knowledge Graph for Explaining Thoracic Pathologies
Appendix A Dataset and Instruction Prompts
For our study, we constructed an instruction-format dataset based on the LLaVA (Li et al. 2024) framework, with specific modifications to the prompts. These modifications were designed to tailor the questions to elicit detailed explanations from the model regarding the reasoning behind the occurrence of specific pathologies observed in the images. The following questions were incorporated into our prompts to guide the model in generating explanations: Which signs show that the patient has pathologies?, Explain why these pathologies are present in the image?, What evidence in the image indicates pathologies?, How can you tell that the patient has pathologies from the image?, What features suggest the presence of pathologies in this image?. These tailored prompts were crucial in enhancing the model’s ability to provide reasoning for the presence of specific pathologies based on the visual evidence present in the images.
Appendix B Implementation Details
Training
The KGR process is powered by the MedCLIP model, which projects the input image into a shared latent space. This projected representation is then used to retrieve corresponding triplets from the KG using the FAISS library (Johnson, Douze, and Jégou 2019), which employs a k-nearest neighbors algorithm to identify the most relevant triplets. These retrieved triplets serve as supplementary information, enriching the input to the language model and improving the model’s ability to generate accurate and contextually relevant explanations.
Training Setup
We trained the LLaVA (Li et al. 2024) model with the integrated KGR module on the MIMIC-NLE dataset. The training process involved fine-tuning the model to maximize the likelihood of generating high-quality NLEs using the additional knowledge provided by the KG.
The training was conducted with LoRA (Hu et al. 2022) (Low-Rank Adaptation) for training efficiency. We used the pre-trained LLaVA (Li et al. 2024) model as the baseline, with a learning rate of 2e-4, and employed the cosine learning rate scheduler. The training was performed over 5 epochs with a batch size of 8 per device and a gradient accumulation of 2 steps. The vision encoder used was CLIP ViT-L (Radford et al. 2021), and the multimodal projector was an MLP with two layers using the GELU activation function.
In addition to our KG-LLaVA framework, we also applied our KGR module to Med-XPT architecture, which utilizes MedCLIP as the vision encoder, a transformer-based projector, and GPT-2 as the language model. We followed a similar structured training approach. We again used the FAISS (Johnson, Douze, and Jégou 2019) library for retrieval based on image features. The training script included several key parameters: the visual encoder (medclip-vit) was used to retrieve captions stored in a JSON file, with the retrieval results incorporated into the input prompts. The attention mechanism within the model was tuned using a cross-attention size of 7, and we enabled the training of the decoder along with the attention mechanism. The training process spanned 15 epochs with a learning rate of 1e-4 and a batch size of 8, utilizing gradient accumulation with one step. The retrieved captions and templates were dynamically integrated into the training pipeline, which facilitated the model’s ability to generate accurate NLEs based on the retrieved knowledge.
Appendix C Impact of Different RAG Methods
@@
Despite the significantly smaller size and less extensive training of the GPT-2 language model compared to LLaVA, Med-XPT—which utilizes GPT-2—achieves competitive performance, particularly when augmented with the KG and NLE-based RAG methods. This suggests that integrating domain-specific knowledge can effectively mitigate the limitations of smaller models, enabling them to generate high-quality explanations.
@@
Notably, when comparing Med-XPT to the DPT (Kayser et al. 2022) framework, which also uses GPT-2 but without the benefit of these advanced RAG methods, we observe a substantial improvement in performance. DPT’s lower scores across the board highlight the critical role that the KG and NLE-based RAG play in enhancing the model’s explanatory capabilities. This further emphasizes that even smaller models, when equipped with the right augmentation techniques, can perform on par with larger, more complex models like LLaVA.
@@
Appendix D Ablation Study on Instruction Prompts
To evaluate the role of instructional prompts in guiding the quality of Natural Language Explanations (NLEs), we conducted an ablation study on 20 randomly selected test instances from the dataset. Our study analyzed the impact of different prompt configurations on BLEU-4, ROUGE-L, and CIDEr scores, which measure precision, recall, and contextual relevance, respectively.
We considered three configurations:
  • Original Prompt (Full Set of Questions): The complete set of instructional questions used in the baseline configuration.
  • Additional Broad Question: Adding the general question, “What patterns in the image are associated with specific pathologies?”
  • No Instructional Questions: All instructional prompts were removed.
The results of the ablation study are summarized in Table 4.
Table 4: Ablation Study Results for Instructional Prompts
Prompt Configuration B-4 R-L CIDEr
Original Prompt (Full Set) 7.56 25.33 82.59
Additional Broad Question 5.24 24.62 63.24
No Instructional Questions 6.70 25.52 70.77
Key Observations
  • Original Prompt (Full Set of Questions): The baseline configuration achieved the highest scores across all metrics (BLEU-4: 7.56, ROUGE-L: 25.33, CIDEr: 82.59). This indicates that a comprehensive set of instructional questions helps produce structured, coherent, and clinically relevant explanations.
  • Additional Broad Question (“What patterns in the image are associated with specific pathologies?”): Adding a broader question reduced BLEU-4 (5.24) and CIDEr (63.24), while ROUGE-L (24.62) remained close to the baseline. This suggests that overly general questions introduce ambiguity, leading to less precise and less targeted explanations.
  • No Instructional Questions: Removing all instructional questions significantly reduced performance, with BLEU-4 at 6.70 and CIDEr at 70.77, though ROUGE-L (25.52) stayed near the baseline. This demonstrates that instructional prompts are crucial for guiding the model towards specific, relevant, and structured outputs.
The ablation study highlights the importance of instructional prompts in enhancing NLE quality. Well-designed prompts focused on specific diagnostic features yield the best performance. Omitting prompts or introducing overly broad questions reduces explanation quality across precision, recall, and contextual relevance. These findings reinforce the role of instructional prompts as a critical component in our framework for generating accurate and coherent explanations.
Appendix E Failure Case Analysis
In this section, we present five failure cases where the KG-LLaVA model demonstrated the lowest performance in terms of CIDEr and ROUGE-L scores. Each case highlights specific challenges faced by the model, such as ambiguity in retrieved triplets, misinterpretation of image features, and difficulty in localizing abnormalities accurately. These insights provide a deeper understanding of the model’s limitations and suggest pathways for improvement.
Table 5: Case 1: Misinterpretation of Opacity Location
Input Image [Uncaptioned image]
Triplets opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions
Question Which signs show that the patient has positive Lung Lesion, positive Lung Opacity, positive Pleural Effusion, positive Pneumonia?
Prediction There is a new opacity in the right lower lobe with a small amount of pleural fluid tracking along the lateral margin of the lung.
Ground Truth (GT) New bibasilar patchy airspace opacities, concerning for aspiration pneumonia.
Table 6: Case 2: Ambiguity in Diagnosing Edema and Effusion
Input Image [Uncaptioned image]
Triplets radiopacity suggestive_of edema; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions
Question Explain why these positive Atelectasis, uncertain Edema, positive Lung Opacity, positive Pleural Effusion are present in the image.
Prediction Bibasilar opacities are likely atelectasis.
Ground Truth (GT) A hazy opacity over the left lung base suggests a layering pleural effusion.
Table 7: Case 3: Inconsistent Interpretation of Pleural Effusion
Input Image [Uncaptioned image]
Triplets radiopacity suggestive_of edema; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions; opacities suggestive_of effusions
Question Explain why these positive Atelectasis, uncertain Edema, positive Lung Opacity, uncertain Pleural Effusion are present in the image.
Prediction Bibasilar opacities, likely atelectasis.
Ground Truth (GT) A small area of hazy opacity at the right costophrenic angle may represent a small layering pleural effusion.
Table 8: Case 4: Limited Context for Multifocal Pneumonia
Input Image [Uncaptioned image]
Triplets line suggestive_of hydropneumothorax; Tubes suggestive_of drain; opacification suggestive_of consolidation; opacification suggestive_of consolidation; opacification suggestive_of consolidation
Question What evidence in the image indicates positive Lung Opacity, positive Pleural Effusion, positive Pneumonia?
Prediction There is a new opacity in the right lower lobe which is concerning for pneumonia.
Ground Truth (GT) Multiple opacifications likely represent multifocal pneumonia, possibly due to aspiration.
Table 9: Case 5: Confusion Between Atelectasis and Effusion
Input Image [Uncaptioned image]
Triplets atelectasis suggestive_of scarring; atelectasis suggestive_of scarring; Tubes suggestive_of drain; opacification suggestive_of collapse; opacification suggestive_of collapse
Question What features suggest the presence of positive Atelectasis, positive Lung Opacity, positive Pleural Effusion in this image?
Prediction There is a left retrocardiac opacity and likely bilateral pleural effusions.
Ground Truth (GT) Substantial bibasilar opacification can be explained by atelectasis.
Case 1 presented in Table 5, the model focused on a right lower lobe opacity, whereas the ground truth describes bibasilar opacities. The repeated and generic triplets (“opacities suggestive of effusions”) provided insufficient contextual specificity. The model’s failure to capture the bilateral nature of the opacities led to an incorrect interpretation of the location, reducing explanation accuracy. This highlights the need for more precise and diverse triplets to improve spatial reasoning.
Case 2 presented in Table 6 shows that the model defaulted to atelectasis as an explanation, likely influenced by the ambiguity in the retrieved triplets, which mention both “radiopacity suggestive of edema” and “opacities suggestive of effusions”. The overlap between findings related to atelectasis and pleural effusion caused the model to miss the correct interpretation. More discriminative triplets distinguishing between these pathologies could mitigate this issue.
Case 3 presented in Table 7 shows that the model leaned towards atelectasis rather than pleural effusion. The lack of specificity in the retrieved triplets led to an incorrect explanation. This case illustrates the model’s difficulty in handling subtle differences between overlapping radiological findings, emphasizing the need for triplets that capture nuanced distinctions.
Case 4 presented in Table 8 shows that the model’s prediction missed the multifocal nature of the pneumonia. The retrieved triplets emphasized “consolidation” but lacked context about the distribution pattern. This case highlights the need for triplets that describe multifocal or distributed findings to guide more accurate explanations.
Case 5 presented in Table 9 shows that the model incorrectly identified bilateral pleural effusions instead of atelectasis. The triplets retrieved (“opacification suggestive of collapse”) did not sufficiently differentiate between collapse and effusion. Enhancing the triplet selection process to provide clearer distinctions between these findings could improve the model’s accuracy.
Summary of Insights
The failure cases reveal several key challenges:
  • Ambiguity in Triplets: Repeated or generic triplets introduce uncertainty, causing the model to default to common diagnoses.
  • Localization Errors: The model struggles to correctly identify the location of abnormalities when multiple findings are present.
  • Lack of Context for Complex Diagnoses: Triplets do not always capture multifocal or distributed patterns, limiting the model’s ability to handle complex cases.
Future improvements can focus on refining the knowledge graph with more specific and context-aware triplets, enhancing spatial reasoning capabilities, and incorporating multi-hop reasoning to improve diagnostic accuracy.
Appendix F Effect of the Number of Retrieved Knowledge Graph Triplets (K) on NLE
@@
Table 10: Effect of the Number of Retrieved Knowledge Graph Triplets (K) on NLE Performance. The evaluation metrics include BLEU-4 (B4), METEOR, ROUGE-L (R-L), and CIDEr.
@@
We conducted an ablation study to investigate the impact of varying the number of retrieved Knowledge Graph (KG) triplets $K$ on the quality of the generated Natural Language Explanations (NLEs). Table 10 presents the evaluation results across BLEU-4 (B4), METEOR, ROUGE-L (R-L), and CIDEr scores.
When retrieving a single triplet ($K=1$), the model achieves strong baseline performance (BLEU-4: 7.3, METEOR: 15.4, ROUGE-L: 25.3, CIDEr: 60.6). This indicates that even a minimal amount of retrieved knowledge can provide sufficient context for generating coherent and relevant explanations.
As $K$ increases to 3 and 5, the performance remains consistent, with minor fluctuations in BLEU-4 (7.1 to 7.2) and METEOR (15.3). The ROUGE-L scores exhibit a slight decline (25.0 to 24.9), suggesting that while additional triplets enrich the context, they do not significantly improve recall-based measures. The CIDEr scores for $K=3$ (58.2) and $K=5$ (58.6) reflect a minor drop compared to $K=1$, indicating that excessive information can introduce noise, potentially reducing the precision of the explanations.
When $K$ is increased to 7, the CIDEr score peaks at 62.2, demonstrating that a higher number of retrieved triplets can enhance contextual relevance and diversity, particularly for complex cases. This improvement in CIDEr suggests that retrieving more knowledge can increase the likelihood of including relevant terms and concepts in the generated explanations. However, the BLEU-4 (7.2) and METEOR (15.1) scores remain stable, indicating that beyond a certain threshold, additional triplets contribute more to contextual richness than to syntactic precision.
The results in Table 10 highlight a trade-off between the quantity of retrieved knowledge and the quality of NLEs. While retrieving more triplets ($K=7$) can improve contextual relevance, as evidenced by the higher CIDEr score, the benefits plateau for precision and recall-based metrics like BLEU-4 and ROUGE-L. Therefore, selecting an optimal $K$ depends on balancing contextual richness with the risk of introducing noise. For our KG-LLaVA model, $K=7$ provides the best overall balance, enhancing the model’s ability to generate informative and contextually relevant explanations.
Table 11: Performance Comparison of Uni-modal and Cross-modal Retrieval on NLE Performance - This table compares the performance of the KG-LLaVA model when using Uni-modal versus Cross-modal retrieval methods for generating Natural Language Explanations (NLEs).
@@
Appendix G Comparison of Uni-modal and Cross-modal Retrieval on NLE Performance
The Cross-modal retrieval method demonstrates a clear advantage over the Uni-modal approach across all evaluation metrics as shown in Table 11, with a particularly notable improvement in CIDEr, where the score increases from 49.9 to 62.2. This substantial gain highlights the effectiveness of integrating both visual and textual modalities in the retrieval process, allowing the model to generate more contextually relevant and accurate Natural Language Explanations (NLEs). The Cross-modal approach enhances the model’s ability to interpret complex medical images, leading to higher-quality explanations that are more aligned with clinical expectations.